# Towards Context-Aware Domain Generalization: Representing Environments with Permutation-Invariant Networks

### 15 Dec 2023 v1

@@ 2.6 Source Component Shift
Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satifsied. Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [46, Chapter 1.9]. The source component shift can be described by the graphical model in Figure 1, where the environment directly affects both the input $\mathbf{X}$ and the outcome $\mathbf{Y}$. Problems that conform to the graph in Figure 1 have two important implications. First, the input distribution changes whenever the environment changes. [^2] In our case, we require a stronger assumption, namely, that the domains are not deducible from a single input, as formalized in Criterion 2.2. Second, the relationship between inputs and outcomes varies with the environment (corresponding to Criterion 2.1). For more details on this kind of distribution shift, we refer the reader to [46, Chapter 1.9]. It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [45, 57], which supplies a proof-of-concept for our approach (see Experiment 1).
@@ 2.7 Detection of Novel Environments
During test time, data could either originate from an environment that corresponds to one of the training environments (but its origins are unknown) or from a previously unseen environment. In the following, we explain how we aim to detect the second case that might result in potential failure cases due to fundamental challenges in extrapolation. Following [41], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$. As a score function, we consider the distance of $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to the $k$-nearest neighbors in the training data in the feature space of the set-encoder. Accordingly, set-representations that elicit a score surpassing a certain threshold are considered to originate from a novel environment. Following the approach in [41], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment. To establish this threshold, we consider the $q$-th percentile of scores obtained from the validation set. We also compare our novel environment detector with the same score function computed solely from singleton features $g(\mathbf{X})$. These results can be quickly previewed in Table 2.
@@ 3.1 Domain Generalization
The aim of domain generalization (DG) is to train models that generalize well under distribution shifts [39, 74]. The DG setting involves access to data from multiple domains during training exploiting the heterogeneity between domains. In this context, the plethora of algorithms aimed at improving robustness has been divided into three categories [60]: (i) algorithms that enhance the data (e.g., via data augmentation or domain synthesis [56, 68]), (ii) algorithms that adapt the learning algorithm (e.g., self-supervised learning [29, 8]), and (iii) algorithms aiming to find robust representations (e.g., via invariance learning [4, 42]). Our approach explores a different avenue exploiting contextual information about the origin of the data via an adaptive environment embedding.
In contrast to domain adaptation (DA) [61], where samples from the test domain are given during training, in DG no knowledge about the test environment is available during training. As a middle-ground between DA and DG, test-time adaptation (TTA) involves the provision of unlabeled samples during test time, enabling further fine-tuning of the model [35]. TTA can be categorized into test-time domain adaptation, test-time batch adaptation (TTBA), and online test-time adaptation (OTTA) [35].
Among these settings, our work aligns most closely with the TTBA scenario concerning the utilization of environment information. In TTBA a pre-trained model is adapted to one or a few inputs [54, 72, 51]. Each adaptation depends on the mini-batch at hand. Similarly, we consider a mini-batch as a set input to deliver contextual information. However, we do not adapt or fine-tune our model at test time, but rather directly extract context information via the set-encoder. This allows us to identify whether extrapolation is necessary, enabling model misspecification detection [50]. Moreover, since we do not need to fine-tune the model at test time, our approach is considerably more efficient.
Finally, [12] assume a setting where inputs from multiple domains are available, but it is unknown which sample belongs to which environment – even during training time. The authors infer potential domain labels that are used for downstream invariant learning, for example, via Invariant Risk Minimization (IRM) [4]. This method improves on baseline models and enables the application of DG methods in the absence of domain labels. Interestingly, it could be combined with our approach when no environment labels are provided during training, and we leave this as an avenue for future research.
@@ 3.2 Learning Permutation-Invariant Representations
Analyzing set-structured data with neural networks has received much theoretical [58, 7, 43] and empirical [69, 33, 70] momentum in recent years. For instance, [70] build on the set transformer architecture [33] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling. The resulting “PICASO” blocks include consecutive multi-head attention stacks with skip connections, which update their template(s) depending on previous template(s) and the particular input set (in contrast to global templates). Differently, [13] propose to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion. The authors also show how to use their criterion for generative and few-shot classification tasks In a somewhat similar vain, [5] investigate a gradient-based optimization method for aggregating set-structured data. Importantly, these methods comprise a pool of algorithms which can be used as the backbone architecture for realizing the set-encoder in our approach.

### 21 Feb 2024 v2

@@ 2.4 Source Component Shift
Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satifsied. Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [Quinonero-Candela et al., 2008]. The source component shift can be described by the graphical model in Figure 1, where the environment directly affects both the input $\mathbf{X}$ and the outcome $\mathbf{Y}$. Problems that conform to the graph in Figure 1 have two important implications. First, the input distribution changes whenever the environment changes. [^1] In our case, we need a stronger assumption: the domains cannot be deduced from a single input, as defined in Criterion 2.2. Second, the relationship between inputs and outcomes varies with the environment (corresponding to Criterion 2.1). For more details on this kind of distribution shift, we refer the reader to Quinonero-Candela et al. [2008]. It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021], which supplies a proof-of-concept for our approach (see Experiment 1). An important point to highlight is that the frequently encountered covariate shift where only $P(\mathbf{X})$ in $P(\mathbf{X},\mathbf{Y})=P(\mathbf{Y}\mid\mathbf{X})P(\mathbf{X})$ varies between environments [Quinonero-Candela et al., 2008], does not conform to the conditions specified in Criterion 2.3. Hence, context-aware models do not provide advantages when compared to standard models under covariate shift (under the assumption all models converged to the optimum).
@@ 2.5 Detection of Novel Environments
During test time, data could either originate from an environment that corresponds to one of the training environments (but its origins are unknown) or from a previously unseen environment. In the following, we explain how we aim to detect the second case that might result in potential failure cases due to fundamental challenges in extrapolation. Following Müller et al. [2023], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$. As a score function, we consider the distance of $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to the $k$-nearest neighbors in the training data in the feature space of the set-encoder. Accordingly, set-representations that elicit a score surpassing a certain threshold are considered to originate from a novel environment.
Following the approach in Müller et al. [2023], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment. To establish this threshold, we consider the $q$-th percentile of scores obtained from the validation set. We also compare our novel environment detector with the same score function computed from singleton features $g(\mathbf{X})$ alone (see Table 1 for a preview).
@@ 3.1 Domain Generalization
The aim of Domain Generalization (DG) is to train models that generalize well under distribution shifts [Muandet et al., 2013, Zhou et al., 2022]. In contrast to domain adaptation (DA) [Wang and Deng, 2018], where samples from the test domain are given during training, in DG no knowledge about the test environment is available during training. Both, DG and DA, involve access to data from multiple domains during training exploiting the heterogeneity between domains. For an overview on DG see for instance Wang et al. [2022], Zhou et al. [2022].
As a middle-ground between DA and DG, test-time adaptation (TTA) [Liang et al., 2023] and marginal transfer learning [Blanchard et al., 2021] involves the provision of unlabeled samples during test time. In TTA, the model is typically fine-tuned according to the provided samples [Liang et al., 2023]. Marginal transfer learning in DG refers to the setting where access to the marginal feature distribution is given during testing and provided to the model via $\frac{1}{n}\sum_{i=1}^{n}\sigma(\mathbf{X}_{i})$ [Blanchard et al., 2021]. In this context, $\sigma$ represents a feature extraction model applied to each $\mathbf{X}_{i}$, and the summation ensures that the operation remains invariant under permutations. Different choices for $\sigma$ are explored in the literature. Specifically, Zhang et al. [2021] utilizes CNNs, Blanchard et al. [2021], Dubey et al. [2021] incorporates kernel embeddings, and Bao and Karaletsos [2023] adopts patch embeddings. While these works highlight the advantages of their methodologies, they fall short in analyzing the exact settings and conditions that contribute to these benefits compared to a standard model. Additionally, they do not address the detection of potential failure cases or the utilization of context embedding for model selection. A recent method proposes to use transformers to exploit previously seen samples instead of learning permutation-invariant embedding [Gupta et al., 2023]
The trade-off between ID performance and OOD performance is well-known [Müller et al., 2022, Magliacane et al., 2018, Zhang et al., 2023]. A method to mitigate this trade-off was proposed in Zhang et al. [2023]. Their approach presupposes knowledge of the data’s originating environment, whereas our goal is to infer it.
@@ 3.2 Learning Permutation-Invariant Representations
Analyzing set-structured data with neural networks has received much theoretical [Wagstaff et al., 2022, Bloem-Reddy and Teh, 2020, Murphy et al., 2018] and empirical [Zaheer et al., 2017, Lee et al., 2019, Zare and Van Nguyen, 2021] momentum in recent years. For instance, Zare and Van Nguyen [2021] build on the set transformer architecture [Lee et al., 2019] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling. Differently, dan Guo et al. [2021] proposes to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion.
Notably, the methods above impose no probabilistic structure on the set-representations, since the latter are mainly used as deterministic features for downstream tasks. In the Bayesian literature, sets represent finitely exchangeable sequences, embodying the core probabilistic structure of most Bayesian models [Orbanz and Roy, 2014]. Typically, hierarchical (aka multi-level) Bayesian models are used to deal with Simpson’s paradox by organizing observations into clusters or levels [Wikle, 2003, Gelman et al., 2013], mirroring the notions of domain or environment in DG. Indeed, hierarchical Bayesian models have been successfully applied in many areas of science, but are typically constrained to linear or generalized linear models (GLMs), prioritizing interpretability over predictive performance.
The machine learning literature is rich with variational methods for learning invariant summary statistics from set-structured data [Edwards and Storkey, 2016, Garnelo et al., 2018, Kim et al., 2019, 2021, Zeng et al., 2022]. Crucially, none of the above methods pursues the concrete goal of our work, which is improving domain generalization performance through set-based environment representations. To achieve this goal, we employ variants of the DeepSet [Zaheer et al., 2017] and SetTransformer [Lee et al., 2019] architectures for our backbone set-encoder throughout all experiments. Interestingly, we observe that our approach is widely robust to the choice of set-encoder architectures in the problems considered.
Figure 2: Experiment 1. Left: Toy dataset that conforms to our theoretical criteria. A detailed description can be found in subsection E.1. Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3. Marginal distributions of the five environments are shown on top. Right: Comparison of environment classification accuracy for different set sizes. Distances between environments refer to the distance between the means of the environments.
3.3 OOD Detection and Selective Classification
Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [Aggarwal and Yu, 2001, Yang et al., 2021, Shen et al., 2021, Han et al., 2022, Yang et al., 2022a]. Flagging out-of-distribution (OOD) instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [Yang et al., 2021]. OOD detection is closely related to inference with a reject option (also termed selective classification) [Geifman and El-Yaniv, 2017, El-Yaniv et al., 2010], which allows classifiers to refrain from predicting ambiguous or novel conditions [Hendrickx et al., 2021]. The reject option has been extensively studied in statistical and machine learning [Hellman, 1970, Fumera and Roli, 2002, Grandvalet et al., 2008, Wegkamp and Yuan, 2012], with early work dating back to the 1950s [Chow, 1957, 1970, Hellman, 1970].
More recently, Müller et al. [2023] explored the utility of selective classification in DG settings. They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent. In this work, we consider a post-hoc score based on the $k$-nearest neighbours to the training set in feature space similar to Sun et al. [2022], which applies to both classification and regression settings. Unlike the approach taken in Müller et al. [2023], where the focus lies on features of individual instances, we consider the set summary provided by the set-encoder. Thus, we can identify novel environments even when singleton inputs lack sufficient information.
4 Experiments
In the following, we explore various aspects of the context-aware model across three different dimensions. First, we show on two datasets that a context-aware model achieves improved performance in the ID as well as the OOD setting compared to a baseline model when the conditions of a source component shift are met. Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model. We also show that novel environment detection can be utilized to avoid failure cases. Third, we demonstrate that the necessary criteria (see subsection 2.3) can be validated empirically, identifying cases where no benefits of the method can be expected. Experimental details can be found in the Appendix. Furthermore, in Appendix B, we conduct an additional experiment resembling Experiment 1, but with high-dimensional inputs
Figure 3: Experiment 1. Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means no improvement is achieved) on a toy example. We also show I (OOD) on OOD data. II depicts the relative improvement of the environment-oracle model compared to the baseline model. III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it. Sampling variation arises from using different seeds to partition the ID data into training, test and validation set.
4.1 Evaluation Approach
Here, $\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})$ denotes a performance measure for our model and similarly for the baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$, both measured on an unseen test set. $\mathcal{R}_{\text{I}}>0$ signifies an advantage attained by the context-aware approach and therefore the fulfillment of Criterion 2.1. In the regression setting, we consider the negative L2-Loss as the performance measure.
To validate Criterion 2.2, we train a contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ utilizing both the set input $\mathbf{S}^{(n)}$ and the target input $\mathbf{X}$ to predict the environment label $E$. Additionally, we train a baseline environment model $f^{E\mid\mathbf{X}}$ aimed at predicting $E$ solely from $\mathbf{X}$. We then compute the relative improvement $\mathcal{R}_{\text{II}}$ [^2] The exact definition for $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ can be found in Appendix D of the contextual environment predictor relative to the baseline environment predictor. $\mathcal{R}_{\text{II}}>0$ indicates that Criterion 2.2 is satisfied. In our experiments, we choose the set size $n$ such that we achieve approximately 100% accuracy for our contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ on ID data.
4.2 Experiment 1: Toy Example
Setup
To set the stage, we consider a dataset shown in Figure 2 and explained in detail in subsection E.1. The dataset includes data from five different environments, defined by distinct Gaussian distributions. Each Gaussian deviates due to its location (i.e. mean vector). The dataset exemplifies Simpson’s paradox, wherein a naive fit of the data (without accounting for environmental factors) would yield a negatively sloped line (see Figure 10). Conversely, using environmental information for prediction enables to capture the true underlying model with positively sloped lines, counteracting the overall negative trend of the mean vectors.
Importantly, the dataset meets our necessary criteria: We cannot infer the origin environment from a single input alone, as indicated by the overlap between the marginal distributions of $X$, obtained by projecting the entire sample onto the $x$-axis in Figure 2. Thus, the setting aligns with Criterion 2.2, and corresponds to Criterion 2.3, since knowing the environment location $\boldsymbol{\mu}^{(j)}$ should improve prediction.
Results
As a first check of Criterion 2.2, we evaluate whether a set input provides additional information about the environment compared to a singleton input. Figure 2 illustrates that additional set input improves the ability to distinguish between environments significantly and the more samples we include, the better the distinction. As expected, a decrease in the distance between environment marginal means necessitates more samples to differentiate between environments. Interestingly, the particular choice of architecture for the permutation-invariant network does not seem to play a significant role for predicting the environment label well, as demonstrated in Appendix I.
Next, we assess the predictive capabilities of the context-aware approach across all possible scenarios of “leave-one-environment-out”. This involves training on all environments except one and treating the excluded environment as a novel OOD scenario. Here, we consider linear models to ensure an optimal inductive bias for the problem (non-linear models achieve similar results, as shown in subsection E.3). We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3. Providing contextual information in the form of a set input increases the performance significantly compared to a baseline model in the ID as well as in the OOD setting (see I and I (OOD) in Figure 3). We also observe a slightly higher relative improvement when the environment label is directly provided (see II) compared to using the output of the set-encoder (see I). This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself. Note that for III we achieve less relative improvement since we consider the accuracy here and not the L2-Loss.
Finally, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 4 for one trained model. Our model captures and utilizes the characteristics of each environment for prediction. In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences. Note that we obtained the best results by considering a class of linear models that aligns with the data-generating process. However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see subsection E.3).
4.3 Experiment 2: Colored MNIST
Setup
The ColoredMNIST dataset [Arjovsky et al., 2019] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (all standard labels $<5$ are assigned to new label 0, and all labels $\geq 5$ are the new label 1). Furthermore, label noise is deliberately added, so only in 75% of all cases, the label can correctly be predicted from the shape. To make things more challenging, the image background can take two colors that are also associated with the image label. In the first environment, the association is 90% and in the second one 80%. Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape. However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy – worse than random.
This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [Arjovsky et al., 2019, Zhang et al., 2023]. For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of diminished accuracy in the first two environments (75% instead of 80% or 90%). In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
[…]
ID OOD
Baseline $\mathbf{84.6}\pm 0.3$ $10.2\pm 0.3$
Invariant $72.8\pm 0.9$ $\mathbf{73.1}\pm 0.2$
Selection (Ours) $\mathbf{84.1}\pm 0.3$ $\mathbf{73.1}\pm 0.2$
Selection (Baseline) $\mathbf{84.0}\pm 0.3$ $14.0\pm 0.4$
Bayes Optimal Classifier $\mathbf{85.0}$ $\mathbf{75.0}$
Table 1: Experiment 2. Mean and standard deviation of accuracy percentages across model types and domain settings over 5 runs. The features extracted by our model allow for improved OOD detection compared to the features of the baseline model. Thus, our model can perform a favorable selection between the baseline model in the ID setting and the invariant model in the OOD setting.
Results
Here, we assume the invariant model to be given (see Appendix G for details), but it could also be obtained by invariant learning, e.g. Invariant Risk Minimization [Arjovsky et al., 2019]. With our novel environment detection approach (see subsection 2.5) we can get the best of both worlds, circumventing the inherent trade-off. When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments. In case we detect the OOD setting, we employ the invariant model. We compare this kind of model selection due to the features $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ inherent to our model versus the features extracted by the baseline model. The results can be found in Table 1. By utilizing the model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data. Evidently, the novel environment detection only works with set summaries. A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 1.
4.4 Experiment 3: Violated Criteria
Setup
We consider the PACS dataset [Li et al., 2017], training our model on the Cartoons, Sketches, and Paintings environments, and assess its performance in the Art environment during testing. The dataset includes images with labels that we intend to predict. Regarding a second classification task, we delve into the OfficeHome dataset [Venkateswara et al., 2017]. In line with the PACS dataset, we approach the classification problem, training across three specific environments, and subsequently evaluating a novel one as an out-of-distribution (OOD) scenario.
(a) Environment Art in PACS dataset. The environment is almost completely inferable from one input sample (Criterion 2.2 not satisfied). Conclusively the context-aware approach does not yield benefits.
Figure 5: Experiment 3. Tell-tale examples where at least one of the necessary criteria is not satisfied and the context-aware approach cannot possibly yield benefits. For experimental details, see Appendix H.
Results
When the criteria are not met, no benefits can be achieved, even in the ID setting. This has been proven in Theorem 2.1 and we demonstrate it here for two scenarios empirically (see Figure 5). We find that Criterion 2.2 is not satisfied on the PACS dataset: As depicted in 4(a), the contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ does not perform better compared to the baseline environment model $f^{E\mid\mathbf{X}}$. Remarkably, a single example is sufficient to infer the source environment, allowing for a 99.7% accuracy in predicting the correct environment from an individual sample (see Appendix H). Also Criterion 2.3 is not satisfied as 4(a) depicts. Since the criteria are not met, we do not achieve any benefit over the baseline model, neither in the ID nor in the OOD setting, as demonstrated in 4(a).
On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is. Results are depicted in 4(b). We observe that the set input offers benefits for predicting the data originating environment corresponding to Criterion 2.3. However, even when providing the target classifier with the environment label (environment-oracle model), we do not achieve an improvement over the baseline model suggesting that Criterion 2.2 is not satisfied. As expected and depicted in 4(b), our method does not yield benefits compared to the baseline model.
4.5 Experiment 4: Failure Case Detection
Setup
Besides unfulfilled criteria, another reason why a context-aware approach might fail to reap benefits is when the distribution shift requires extrapolation. This might be unattainable by the model, making the inclusion of a reject option beneficial. Using the BikeSharing dataset [Fanaee-T, 2013], we demonstrate that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary. For this dataset we consider the task of predicting the number of bikes rented across the day based on weather data. Here we explore the scenario where we train on all seasons except winter. Details about the dataset, pre-processing steps and other scenarios can be found in Appendix J.
Results
In Table 5 we demonstrate that the context-aware approach is slightly superior compared to the baseline model in the ID settings. However, both the baseline and the context-aware approach experience performance degradation in the novel winter environment. To detect the novel environment and, consequentially, potential failure cases, we compute the score as suggested in subsection 2.5 and evaluate how well it distinguishes between ID versus OOD environments. We designate an independent ID test set and use the environment excluded during training (here winter) as the OOD set for evaluation. The area under the ROC-curve (AUROC) in Table 5 demonstrates that the score based on the permutation-invariant embedding allows for a perfect detection of the novel environment, whereas the standard approach fails as expected.
Table 2: Experiment 4. Performance comparison between our model and the baseline for the winter domain in the BikeSharing dataset. We compare the performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC). Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case. We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets. Other domains can be found in Appendix J.
5 Conclusions
In this work, we aimed to advance the theoretical understanding of marginal transfer learning in domain generalization. Thus, we formulated criteria that are necessary for the approach to yield benefits and are also easy to verify empirically. Moreover, we pinpoint the source component shift as a scenario where context-aware models can offer advantages, enabling the identification of favorable scenarios. Additionally, we illustrate the capability to identify novel environments, thereby allowing for the detection of potential failure cases. This, in turn, also supports model selection by choosing the most predictive on ID data and the most robust model on OOD data.
Acknowledgments
Appendix A Algorithm
Data: Samples from the joint distribution $p(\mathbf{X},\mathbf{Y},E)$
for $i=1,\dots,k$ do
       Sample mini-batch =$\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
       for $j=1,\dots,m$ do
             Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
             Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
       end for
      Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
end for
Output : Trained context-aware model $f_{\boldsymbol{\theta}}$
A.1 Model Overview
For a detailed overview of the models used to verify the criteria, see Table 3.